A car passes through a gasoline station at an average speed of 36 kph. After 10 seconds another car passes at the same station with an average speed of 72 kph. How long will the second car overtake the first car?
Answer by lwsshak3(11628)

A car passes through a gasoline station at an average speed of 36 kph. After 10 seconds another car passes at the same station with an average speed of 72 kph. How long will the second car overtake the first car?
***
let x=how long second car will overtake first car. (also, travel time of second car)
(x+10 sec)=travel time of first car
distance=travel time *speed(equal for both cars)
hr=3600 sec
10sec=(10/3600)=1/360 hr
..
72x=36(x+1/360)
72x=36x+0.1
36x=0.1
x=0.1/36 hr = .1/36*3600=10sec
how long second car will overtake first car=10 sec